# Kinetic Fragmentation of the QCD Axion on the Lattice

### v1

@@ 1 Introduction
Many alternative scenarios to the vacuum misalignment mechanism have been conceived such as a large [32, 33, 34, 35, 36] or small [37, 38, 39, 40] initial misalignment angle, parametric resonance [41, 42, 43], frictional misalignment [44], trapped misalignment [45, 46], acoustic misalignment [47], bubble misalignment [48], varying axion decay constant [49], curvature induced dark matter [50] interaction with monopoles [51, 52], level crossings [53, 54], modifications in the cosmological history of the universe [55, 56, 13, 57, 58, 59, 60, 61, 62, 63, 64] and finally kinetic misalignment [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80]. The large variety of plausible mechanisms for axion dark matter generation listed above lead to different predictions for the present day axion abundance compared to the standard vacuum misalignment scenario. As a result, it is important to probe experimentally the widest possible parameter space in the $(m_{0},f)$ plane and not just the one associated with the QCD axion or the vacuum misalignment mechanism alone.
@@
Soon after the introduction of kinetic misalignment, it was pointed out that it is typically accompanied by axion fragmentation [77, 84, 85]. The combined scenario that includes the two aforementioned effects is sometimes referred to as “kinetic fragmentation” [79, 80]. This occurs because as the axion zero mode rolls over potential maxima, or oscillates around a minimum with large amplitude, the axion perturbations experience tachyonic instabilities and/or parameteric resonance which exponentially enhances their occupation number. These enhanced perturbations then backreact on the axion, stripping away its energy. Fragmentation occurs when the energy in perturbations is much greater than the energy of the zero mode. Reference [79] is the most comprehensive reference on the effects of fragmentation. The effects of fragmentation, at least for the purpose of axion dark matter generation, have so far been studied analytically or numerically using linear perturbation theory and homogeneous backreaction. The picture that emerges from this analysis is that fragmentation alters the final dark matter abundance compared to the standard kinetic misalignment by an ${\cal O}(1)$ factor. Additionally, due to the highly peaked production of axion perturbations, these perturbations have significant consequences for dark matter mini-halo formation at late times [80]. This is particularly promising as it provides an experimental handle for falsifying this mechanism. Finally, the process of fragmentation can lead to production of gravitational waves whose amplitude can be greater than the gravitational wave background produced during inflation [79]. While the scenario above is very appealing, it remains unclear how including non-perturbative effects and inhomogeneous backreaction may alter these predictions.
@@
Studies of the axion-like particle (ALP) dark matter paradigm using non-perturbative lattice techniques have traditionally focused on the post-inflationary PQ symmetry breaking scenario, in which axions are predominantly produced via topological defects such as strings and domain walls (see [87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101] for a collection of recent works). However, the application of such lattice methods to the study of axion fragmentation in the pre-inflationary scenario is an emerging area of interest, motivated by a variety of recent proposals. These include studies of the relaxion mechanism [85], the effect of monodromy [102, 84] or non-periodic potentials [103], and models involving couplings to hidden dark gauge sectors [104, 105].
The aim of the present work is to perform the first comprehensive fully non-perturbative classical lattice simulation of kinetic fragmentation in the context of axion dark matter. To this end, we limit our analysis to the QCD axion as that is the most well motivated scenario from the perspective of the standard model, and also the most compelling parameter space within reach of axion haloscopes. The parameter space we studied is highlighted in blue in Fig. 1. We narrowed our analysis to this range because for smaller masses, the QCD axion does not fragment and therefore the evolution is well captured by a linear approximation. One the other hand, the upper limit is due to limits in our computational resources. Our goal is to confront the approximate results of linear theory and homogeneous backreaction against the non-perturbative lattice simulations and get a quantitative grasp of how non-perturbative dynamics change the conclusions summarized in the previous paragraph. Our lattice simulations were performed using an adapted version of ${\mathcal{C}}$${\mathcal{L}}$attice [106, 107].
@@ 2.1 Axion dark matter abundance and trapping temperature
$m^{2}(T)=m_{0}^{2}\,\times\,\begin{cases}(T/T_{\rm c})^{-\gamma}&\text{for}\;\;T>T_{\rm c}\\ 1&\text{for}\;\;T\leq T_{\rm c}\,,\end{cases}$ (2.2)
@@ 2.2 QCD axion parametrization and initial conditions
$n$ $f\ (\text{GeV})$ $m_{0}\ (\text{eV})$ $m_{*}\ (\text{eV})$
@@
400 $2.5288\cdot 10^{9}$ $2.2601\cdot 10^{-3}$ $6.6291\cdot 10^{-7}$
@@
We use the subscript “$\rm ini$” to denote variables evaluated at the start of our simulation. The scale factor in radiation domination is $a=(t/t_{\rm ini})^{1/2}=(T_{\rm ini}/T)$ which implies that we are disregarding the change of the effective degrees of freedom during the evolution, and employ their values at the moment of trapping. Our initial conditions are set at an early time when the kinetic energy of the axion is much greater than the potential energy. Let us define a parameter as the ratio between the initial kinetic and potential energies
@@ 3.2 Fully non-perturbative lattice simulations
For the non-perturbative analysis of the dynamics we use lattice simulation. For this end, we employ the basic toolkit of the ${\mathcal{C}}$${\mathcal{L}}$attice library [107] for scalar fields, where the equations of motion are evolved in grids with $N$ points per dimension, $\delta x$ spatial lattice spacing, and with periodic boundary conditions. The background geometry is fixed and assumed to expand at a constant rate, corresponding to radiation domination. In this section, we briefly outline the key aspects of the numerical setup and refer the reader to [106] for further details.
We use the conformal time, $d\tau/d=a^{-1}dt/d=(...)^{\prime}$, version of Eq. (2.7) redefined as
@@
$\begin{split}(\pi_{\theta})_{+\frac{\hat{0}}{2}}&=\pi_{\theta}+\frac{\delta\tau}{2}\mathcal{K}[\theta,a,H]\;,\\ \theta_{+\hat{0}}&=\theta+\delta\tau(\pi_{\theta})_{+\frac{\hat{0}}{2}}a^{-2}_{+\frac{\hat{0}}{2}}\;,\\ (\pi_{\theta})_{+\hat{0}}&=(\pi_{\theta})_{+\frac{\hat{0}}{2}}+\frac{\delta\tau}{2}\mathcal{K}[\theta_{+\hat{0}},a_{+\hat{0}},H_{+\hat{0}}]\;,\end{split}$ (3.8)
@@
As customary in canonical scalar theories, the scalar field $\theta(\mathbf{n})$ and its conjugate momentum $\pi_{\theta}(\mathbf{n})$ are both defined on the sites $\mathbf{n}$ of the lattice, while forward and backward spatial derivatives, $\Delta^{\pm}_{i}\theta(\mathbf{n}+\hat{\imath}/2)=[\pm\theta(\mathbf{n}\pm\hat{\imath})\mp\theta(\mathbf{n}))]/\delta x$, are implemented halfway between neighbour sites. The subscript $+\hat{0}$ denotes a shift of the discrete field by one time step $\delta\tau$. Note that we employ the VV2 scheme in Eq. (3.8), which requires a $+\hat{0}/2$ half-step update of $\pi_{\theta}$ to evolve the system.
@@ 4.2 Comparison of the spectra
starting from the beginning of our simulation $a_{\rm ini}=1$ until the end $a_{\rm end}=2.5$ with a step $\Delta a=0.1$. On the lattice we use the Type II-Version 1 power spectrum, as defined in [110]. We select $n=500$ as a typical representative value and superimpose the spectra of the homogeneous backreaction (in dashed lines) with the ones from the lattice (solid lines). Our results are displayed in Fig. 4. The horizontal axis of the figure has been chosen to be $k/(a\;m(T))$ strategically so that $k/(a\;m(T))=1$ (displayed as a grey, dashed, vertical line) indicates the transition below which modes are non-relativistic. The dispersion relation of the perturbations can be read off Eq. (3.1),
@@ 5.3 Dilution factor compensation algorithm
[…]
@@ Appendix A Initial conditions for the perturbations
[…]
@@
[…]
@@
Where $\Theta^{\prime}$ and $a\,H$ can be evaluated at any moment after $x_{\rm early}$. Since we are evaluating the code initial conditions at a moment when all the modes are well within the horizon, one should in principle derive a more general solution that considers the effects of the rest of the terms beyond the homogeneous one which become important at subhorizon scales. Such contributions however only affect the initial conditions up to ${\cal{O}}(1)$ corrections so in practice keeping only Eq. (A.10) yields a sufficient answer up to the sensitivity required in our simulation.

### v2

@@ 1 Introduction
Many alternative scenarios to the vacuum misalignment mechanism have been conceived such as a large [32, 33, 34, 35, 36] or small [37, 38, 39, 40] initial misalignment angle, parametric resonance [41, 42, 43], frictional misalignment [44], trapped misalignment [45, 46, 47], acoustic misalignment [48], bubble misalignment [49], varying axion decay constant [50], curvature induced dark matter [51] interaction with monopoles [52, 53], level crossings [54, 55], modifications in the cosmological history of the universe [56, 57, 58, 13, 59, 60, 61, 62, 63, 64, 65, 66] and finally kinetic misalignment [67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82]. The large variety of plausible mechanisms for axion dark matter generation listed above lead to different predictions for the present day axion abundance compared to the standard vacuum misalignment scenario. As a result, it is important to probe experimentally the widest possible parameter space in the $(m_{0},f)$ plane and not just the one associated with the QCD axion or the vacuum misalignment mechanism alone.
@@
Soon after the introduction of kinetic misalignment, it was pointed out that it is typically accompanied by axion fragmentation [79, 86, 87]. The combined scenario that includes the two aforementioned effects is sometimes referred to as “kinetic fragmentation” [81, 82]. This occurs because as the axion zero mode rolls over potential maxima, or oscillates around a minimum with large amplitude, the axion perturbations experience tachyonic instabilities and/or parameteric resonance which exponentially enhances their occupation number. These enhanced perturbations then backreact on the axion, stripping away its energy. Fragmentation occurs when the energy in perturbations is much greater than the energy of the zero mode. Reference [81] is the most comprehensive reference on the effects of fragmentation. The effects of fragmentation, at least for the purpose of axion dark matter generation, have so far been studied analytically or numerically using linear perturbation theory and homogeneous backreaction. The picture that emerges from this analysis is that fragmentation alters the final dark matter abundance compared to the standard kinetic misalignment by an ${\cal O}(1)$ factor. Additionally, due to the highly peaked production of axion perturbations, these perturbations have significant consequences for dark matter mini-halo formation at late times [88, 82]. This is particularly promising as it provides an experimental handle for falsifying this mechanism. Finally, the process of fragmentation can lead to production of gravitational waves whose amplitude can be greater than the gravitational wave background produced during inflation [81]. While the scenario above is very appealing, it remains unclear how including non-perturbative effects and inhomogeneous backreaction may alter these predictions.
@@
Studies of the axion-like particle (ALP) dark matter paradigm using non-perturbative lattice techniques have traditionally focused on the post-inflationary PQ symmetry breaking scenario, in which axions are predominantly produced via topological defects such as strings and domain walls (see [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104] for a collection of recent works). However, the application of such lattice methods to the study of axion fragmentation in the pre-inflationary scenario is an emerging area of interest, motivated by a variety of recent proposals. These include studies of the relaxion mechanism [87], the effect of monodromy [105, 86] or non-periodic potentials [106], and models involving couplings to hidden dark gauge sectors [107, 108] and models involving multiple dark sectors [109].
[…]
@@ 2.1 Axion dark matter abundance and trapping temperature
[…]
@@ 2.2 QCD axion parametrization and initial conditions
$n$ f(GeV)f\ ($\text${GeV}) $m_{0}\ (\text{eV})$ $m_{*}\ (\text{eV})$
@@
400 $5288\cdot 10^{9} 2.2$$2601\cdot 10^{-3}$ $6.6291\cdot 10^{-7}$
@@
We use the subscript $\rm ini$” to denote variables evaluated at the start of our simulation. The scale factor in radiation domination is $a=(t/t_{\rm ini})^{1/2}=(T_{\rm ini}/T)$ which implies that we are disregarding the change of the effective degrees of freedom during the evolution, and employ their values at the moment of trapping. Our initial conditions are set at an early time when the kinetic energy of the axion is much greater than the potential energy. Let us define a parameter as the ratio between the initial kinetic and potential energies
@@ 3.2 Fully non-perturbative lattice simulations
[…]
@@
[…]
@@
As customary in canonical scalar theories, the scalar field $\theta(\mathbf{n})$ and its conjugate momentum $\pi_{\theta}(\mathbf{n})$ are both defined on the sites $\mathbf{n}$ of the lattice, while forward and backward spatial derivatives, $\Delta^{\pm}_{i}\theta(\mathbf{n}+\hat{\imath}/2)=[\pm\theta(\mathbf{n}\pm\hat{\imath})\mp\theta(\mathbf{n}))]/\delta x$, are implemented halfway between neighbour sites. The subscript $\hat{0}$ denotes a shift of the discrete field by one time step $\delta\tau$. Note that we employ the VV2 scheme in Eq. (3.8), which requires a $\hat{0}/2$ half-step update of $\pi_{\theta}$ to evolve the system.
@@ 4.2 Comparison of the spectra
starting from the beginning of our simulation $a_{\rm ini}=1$ until the end $a_{\rm end}=2.5$ with a step $\Delta a=0.1$. On the lattice we use the Type II-Version 1 power spectrum, as defined in [114]. We select $n=500$ as a typical representative value and superimpose the spectra of the homogeneous backreaction (in dashed lines) with the ones from the lattice (solid lines). Our results are displayed in Fig. 4. The horizontal axis of the figure has been chosen to be k/(am(T))k/(a\;m(T)) strategically so that k/(am(T))=1k/(a\;m(T))=1 (displayed as a grey, dashed, vertical line) indicates the transition below which modes are non-relativistic. The dispersion relation of the perturbations can be read off Eq. (3.1),
@@ 5.3 Dilution factor compensation algorithm
where ${\cal Z}_{{{\rm test},\,0}}=1$. One also must exercise caution in selecting the value of $n$ at every iteration. In the first iteration the value of $n$ is arbitrary and it fixes the mass and axion decay constant $(m_{0},f)$. In order to maintain the same values in the subsequent iterations, one ought to find the corresponding $n$ value, which leaves the mass and axion decay constant invariant in light of the new trapping temperature, which can be done using Eqs. (2.6) and (2.2).
@@ Appendix A Initial conditions for the perturbations
$\theta^{\rm par}_{k}=\theta_{s}(x)\int^{x}dx^{\prime}\frac{\theta_{c}(x^{\prime}){\cal F}(x^{\prime})}{W\left[\theta_{c}(x^{\prime}),\theta_{s}(x^{\prime})\right]}-\theta_{c}(x)\int^{x}dx^{\prime}\frac{\theta_{s}(x^{\prime}){\cal F}(x^{\prime})}{W\left[\theta_{c}(x^{\prime}),\theta_{s}(x^{\prime})\right]}$ (A.5)
@@
${\cal I}_{c,s}(x)=\int^{x}dx^{\prime}\frac{\theta_{c,s}(x^{\prime}){\cal F}(x^{\prime})}{W\left[\theta_{c}(x^{\prime}),\theta_{s}(x^{\prime})\right]}\;.$ (A.8)
@@
Where $\Theta^{\prime}$ and aHa\,H can be evaluated at any moment after $x_{\rm early}$. Since we are evaluating the code initial conditions at a moment when all the modes are well within the horizon, one should in principle derive a more general solution that considers the effects of the rest of the terms beyond the homogeneous one which become important at subhorizon scales. Such contributions however only affect the initial conditions up to ${\cal{O}}(1)$ corrections so in practice keeping only Eq. (A.10) yields a sufficient answer up to the sensitivity required in our simulation.